Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Basic Dice Probability Question

Expert replies
by hotwann » Mon Aug 27, 2012 4:06 pm
Hi - I had question regarding probabilities with a six-sided die.

The MGMAT Number Properties study guide has the following example question:
What is the probability that, on three rolls of a number cube with faces numbered 1 to 6, at least one of the rolls will be a 6?
So, after reading, I understand that you could calculate this by finding that for each of the three rolls, there is a 5/6 probability that the die will not yield a 6. Or, you can do 1/6+5/6*1/6+5/6*5/6*1/6 to add the probabilities of each of the three scenarios.

So, for some reason, my dumb brain naturally wants to solve this problem by just saying "1/6 + 1/6 + 1/6."

I understand it is not the answer to the question, given that it doesn't not account for the three scenarios correctly. But conceptually, what would 1/6 + 1/6 + 1/6 represent? Is there a way to put that into words? That might help me understand better why it's wrong.

Thanks
Join the discussion
Source: — Problem Solving |

by SmartAssJun » Mon Aug 27, 2012 4:59 pm
hotwann wrote:Hi - I had question regarding probabilities with a six-sided die.

The MGMAT Number Properties study guide has the following example question:
What is the probability that, on three rolls of a number cube with faces numbered 1 to 6, at least one of the rolls will be a 6?
So, after reading, I understand that you could calculate this by finding that for each of the three rolls, there is a 5/6 probability that the die will not yield a 6. Or, you can do 1/6+5/6*1/6+5/6*5/6*1/6 to add the probabilities of each of the three scenarios.

So, for some reason, my dumb brain naturally wants to solve this problem by just saying "1/6 + 1/6 + 1/6."

I understand it is not the answer to the question, given that it doesn't not account for the three scenarios correctly. But conceptually, what would 1/6 + 1/6 + 1/6 represent? Is there a way to put that into words? That might help me understand better why it's wrong.

Thanks
First let's get the probability that none of the rolls are 6.
The probablility of each roll of the dice not being 6 is 5/6
So the probabilty of none of the rolls being 6 should be (5/6)^3
which is 125/216. That's what the probability of that is.
Join the discussion

by SmartAssJun » Mon Aug 27, 2012 5:03 pm
SmartAssJun wrote:
hotwann wrote:Hi - I had question regarding probabilities with a six-sided die.

The MGMAT Number Properties study guide has the following example question:
What is the probability that, on three rolls of a number cube with faces numbered 1 to 6, at least one of the rolls will be a 6?
So, after reading, I understand that you could calculate this by finding that for each of the three rolls, there is a 5/6 probability that the die will not yield a 6. Or, you can do 1/6+5/6*1/6+5/6*5/6*1/6 to add the probabilities of each of the three scenarios.

So, for some reason, my dumb brain naturally wants to solve this problem by just saying "1/6 + 1/6 + 1/6."

I understand it is not the answer to the question, given that it doesn't not account for the three scenarios correctly. But conceptually, what would 1/6 + 1/6 + 1/6 represent? Is there a way to put that into words? That might help me understand better why it's wrong.

Thanks
First let's get the probability that none of the rolls are 6.
The probablility of each roll of the dice not being 6 is 5/6
So the probabilty of none of the rolls being 6 should be (5/6)^3
which is 125/216. So the probability that at least one of the rolls is 6
is 1-(125/216) = 91/216, that's how possibily you'd get 6 at least once out of
three rolls.
Join the discussion

by mehaksal » Tue Aug 28, 2012 3:05 am
shouldn't the answer to this question be :-
1/6*5/6*5/6 + 1/6*/1/6*5/6 + 1/6*1/6*1/6
Join the discussion

by Ian Stewart » Tue Aug 28, 2012 5:49 pm
hotwann wrote: So, for some reason, my dumb brain naturally wants to solve this problem by just saying "1/6 + 1/6 + 1/6."

I understand it is not the answer to the question, given that it doesn't not account for the three scenarios correctly. But conceptually, what would 1/6 + 1/6 + 1/6 represent? Is there a way to put that into words? That might help me understand better why it's wrong.
In probability (and more generally in counting problems, for the same reasons), we add when we have different cases. So if you were asked "what is the probability when you roll a die once that you get a 1, 2 or 5?" then one (long) way to answer the problem is to break it into cases. We can get the result we want in one of three distinct ways: we roll a 1 *or* we roll a 2 *or* we roll a 5. We can add the probability of each case to get the answer: 1/6 + 1/6 + 1/6 = 3/6. So that's the type of situation where you might arrive at the type of answer you wanted to write down here - where there are three distinct ways to get the result the question asks for, and each has a probability of 1/6.

Notice that we do this kind of addition in daily life all the time too - if you know the probability of rain is 40% and the probability of snow is 15%, then (assuming it can't both rain and snow) you would probably quite naturally say that the probability of rain or snow would be 55%, and that would be perfectly correct.

Now, you multiply in probability (and, more generally, in counting) in a completely different situation: when you need a sequence of results to all happen. So if you roll a die three times, and want to know the probability you get a '6' on each roll, you would multiply the probability that happens each time: 1/6 * 1/6 * 1/6 = 1/6^3. So if you ever see a probability problem where you're rolling a die three times, or flipping a coin three times, or whatever, you're usually going to end up multiplying at some point because you're going to have to consider a sequence of three outcomes.

And, of course in more complicated questions, the two situations described above can be combined, so you may need to both add and multiply in the same question - e.g. 'if you flip a coin 3 times, what's the probability you get either 3 heads or 3 tails?'. Then you have two cases (all heads, or all tails), but to find the probability of each case, we'll need to multiply.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion